on the spectrum , label the major absorption peaks with the structrual fragment that causes the absorption. on the answer sheet that follows, fill the requested infomation so that you assemble all the structrual information that can be obtained from the spectra. Then using the informatioon, propose the structure and identity of the compund. A logical analysis of the data to suposrt the identificaion should be written in the "combined analysis" section. You can confirm the identity from literature or on-line sources.

%C 84.8; %H 7.1 4, UV: λmax 217 nm, log e = 3.8 100 80 O 40 20- M+ 198 o-tm 20 40 60 80 10o 120 140 160 180 200 m/z
on the spectrum , label the major absorption peaks
on the spectrum , label the major absorption peaks
Answer #1

From the given data

Elemental analysis

mols of C = 84.8/12 = 7.03

mols of H = 7.1/1 = 7.1

mols of O = 8.1/16 = 0.50

divide by smallest fraction,

C = 7.03/0.50 = 14

H = 7.1/0.50 = 14

H = 0.50/0.50 = 1

Empirical formula = C14H14O

empirical formula = molecular formula

UV analysis

lambda-max = 217 nm : we have aromatic system

IR analysis

3100 cm-1 for alkynic linkage

3000-3050 cm-1 for sp2 =C-H stretch

2850-2950 cm-1 for aliphatic C-H stretch

1500-1600 cm-1 for C=C stretch

Mass analysis

M+ 198

m/z 108 formed by loss of Ph and -CCH fragments

m/z 92 formed by loso of oxygen from 108

Molecular formula

C14H14O